OpenStudy (anonymous):

evaluate the definite integral [0,1] xe^(-x^2)dx and please show steps

OpenStudy (anonymous):

xe^(-x^2)dx = -(1/2) *-2xe^(-x^2)dx substitute with u=-x^2 -1/2 Int[e^u,du] = -(1/2) e^u = -(1/2) * e^(-x^2)

OpenStudy (anonymous):

Evaluate (-(1/2) * e^(-1^2)) - (-(1/2) * e^((-0)^2)) = -1/(2e) + 1/2